A rectangular corral built with 190 meters of fencing that includes four separate grazing areas lies on a 962 meters squared rectangular plot of land. The fence that divides the grazing areas is parallel to two of the sides of the corral. If there is a 1 meter strip of land between each side of the corral and the property line, find the dimensions fo the corral. There may be more than one answer.
Answer by scott8148(6628)

the perimeter (2W+2L) of the corral plus 3 interior fences (3W) is 190, so 5W+2L=190 ... let x=length, so (190-2x)/5=width

the area of the corral (LW) plus the 1 m border is 962 ... (x+2)((190-2x+10)/5)=962

(x+2)(200-2x)=5*962 ... 200x-2x^2+400-4x=4810 ... x^2-98x+2205=0

use quadratic formula to find x